H1. An elephant named Kevin throws a 4 kg ball from 2.6m off the ground. What is the speed of each ball as it hits the ground if a) Ball a is dropped from rest. b) Ball b is thrown straight upward with an initial velocity of 6.0 m/s. c) Ball c is thrown straight downward with an initial velocity of 6.0 m/s. d) Ball d is thrown at an angle of 40 above the horizontal with an initial velocity of 6.0 m/s.
Answer #1

פ Bal, a is da.rr여 トー^..t V E tid nals pee V= um/s Am4 we stth@ v= ? S2.6m 36 +509 V= 9.32 mls v =? S-2. 9 400 - t6.Cos 40 =